Prepare Job-Order Cost Sheets, Predetermined Overhead Rate, Ending Balance of WIP, Finished Goods, and COGS

At the beginning of June, Rhone Company had two jobs in process, Job 44 and Job 45, with the following accumulated cost information:

Job 44 Job 45
Direct materials $5,300 $1,300
Direct labor 1,400 3,100
Applied overhead 980 2,170
Balance, June 1 $7,680 $6,570

During June, two more jobs (46 and 47) were started. The following direct materials and direct labor costs were added to the four jobs during the month of June:

Job 44 Job 45 Job 46 Job 47
Direct materials $2,400 $7,160 $1,850 $1,750
Direct labor 740 6,420 1,000 520

At the end of June, Jobs 44, 45, and 47 were completed. Only Job 45 was sold. On June 1, the balance in Finished Goods was zero.

Required:
1. Calculate the overhead rate based on direct labor cost. Round to three decimal places.
2. Prepare a brief job-order cost sheet for the four jobs. Show the balance as of June 1 as well as direct materials and direct labor added in June. Apply overhead to the four jobs for the month of June, and show the ending balances.
3. Calculate the ending balances of Work in Process and Finished Goods as of June 30.
4. Calculate the Cost of Goods Sold for June.

Refer to the list below for the exact wording of text items within your job-order cost sheets.

Amount Descriptions
Applied overhead
Beginning balance, June 1
Direct labor
Direct materials
Total, June 30

Please show how you got your answers if necessary
Answer #1

1) Overhead rate = applied overhead/Direct labor cost = 980/1400 = 70%

2) Job Cost sheet

Job 44 Job 45 Job 46 Job 47
Beginning balance 7680 6570 0 0
direct material 2400 7160 1850 1750
Direct labor 740 6420 1000 520
Applied overhead 518 4494 700 364
Total 11338 24644 3550 2634

3) Ending work in process = Job 46 = 3550

Finished goods inventory = Job 44+Job 47 = 11338+2634 = 13972

4) Cost of goods sold = Job 45 = 24644